OpenStudy (anonymous):

Product of (2x-2y)(x+2y)

OpenStudy (anonymous):

By distributive property \[\Large (2x-2y)(x+2y)=(2x)(x)+(2x)(2y)+(-2y)(x)+(-2y)(2y)\] \[\Large= 2x ^{2}+4xy-2xy-4y ^{2}\] \[\Large=2x ^{2}+2xy-4y ^{2}=2(x-y)(x+2y)\]

OpenStudy (anonymous):

use FOIL. F-first O-outer I-inner L-last.

OpenStudy (anonymous):

once you reduce it down or simplify- The answer is x^2 + xy - 2y^2.

OpenStudy (anonymous):

(2x−2y)(x+2y) =(2x+−2y)(x+2y) =(2x)(x)+(2x)(2y)+(−2y)(x)+(−2y)(2y) =2x^2+4xy−2xy−4y^2 =2x^2+2xy−4y^2 Just use FOIL.